# MetaMixer Is All You Need

### v1

@@ 3.3 Fast-Forward Network as a General Backbone
Given our proposed FFNified attention and off-the-shelf ConvNeXt block as primary building mixers, we introduce Fast-Forward Network (FFNet), a new network family that demonstrates exceptional performance across a wide range of tasks with task-related modifications. The basic block of our FFNet is illustrated in Fig. 3 (c). Moreover, our model consists solely of hardware-friendly operations, ensuring high efficiency across various devices (GPU and Mobile device).
@@
4 Experimental Results
To demonstrate the versatility of our proposed method, we compare our Fast-Forward Network (FFNet) against the state-of-the-art baselines on diverse tasks such as image classification, object detection and instance segmentation, 2D semantic segmentation, super-resolution, 3D semantic segmentation, and long-term time series forecasting. Additional results can be found in the Appendix.
@@ 4 Experimental Results
Speed Measurement. Following the same protocol as [93, 48], mobile latency is measured on iPhone 12 (iOS 16.5) using models exported by CoreML tools, while GPU throughput is measured with an NVIDIA RTX 4090 GPU. For throughput measurements on the GPU, we compile the traced model with TensorRT (v8.6.1.post1). Unless otherwise specified, all methods are evaluated using their respective input sizes and a batch size of 1, following the aforementioned procedure.
4.1 Image Classification
@@
Table 1: ImageNet-1K classification results. We compare our models with state-of-the-art methods on ImageNet classification.“-” denotes that model could not be reliably exported by either TensorRT or Core ML Tools. The throughput/latency is measured with warmup and averaged over multiple runs.
Model Image Param FLOPs GPU Mobile Top-1 Size Throu. Latency Acc. (Px) (M) (G) (Img/Sec) (ms) (%) EfficientNet-B1 [85] $240^{2}$ 7.8 0.7 854 1.8 79.1 FastViT-SA12 [93] $256^{2}$ 10.9 1.9 731 2.0 80.6 Swin-T [59] $224^{2}$ 29.0 4.5 - 24.8 81.3 FFNet-1 $256^{2}$ 13.7 2.9 1113 1.8 81.3 DeiT3-S [90] $224^{2}$ 22.0 4.6 243 20.8 81.4 CoAtNet-0 [19] $224^{2}$ 25.0 4.2 - 8.2 81.6 CMT-XS [32] $192^{2}$ 15.2 1.5 424 3.9 81.8 ResNetV2-101 [100] $224^{2}$ 44.5 7.8 449 3.6 82.0 ConvNeXt-T [60] $224^{2}$ 29.0 4.5 789 3.7 82.1 PoolFormer-M48 [110] $224^{2}$ 73.5 11.6 - 7.7 82.5 FastViT-SA24 [93] $256^{2}$ 20.6 3.8 439 3.0 82.6 HorNet-T7×7 [74] $224^{2}$ 22.0 4.0 428 3.4 82.8 EfficientNet-B4 [85] $380^{2}$ 19.3 4.2 572 5.2 82.9 Swin-S [59] $224^{2}$ 50.0 8.7 - 29.3 83.0 FFNet-2 $256^{2}$ 26.9 6.0 674 3.1 82.9 CoAtNet-1 [19] $224^{2}$ 42.2 8.4 - 17.0 83.3 MOAT-0 [107] $224^{2}$ 27.8 5.7 - 5.1 83.3 iFormer-S [81] $224^{2}$ 20.0 4.8 468 27.3 83.4 Swin-B [59] $224^{2}$ 87.8 15.4 - 33.5 83.5 CMT-S [32] $224^{2}$ 25.1 4.0 379 11.3 83.5 NFNet-F0 [4] $256^{2}$ 71.5 12.4 548 6.0 83.6 EfficientNet-B5 [85] $456^{2}$ 30.0 9.9 401 11.0 83.6 HorNet-S7×7 [74] $224^{2}$ 49.5 8.8 331 5.4 83.8 ConvNeXt-B [60] $224^{2}$ 88.6 15.4 399 8.1 83.8 EfficientNetV2-S[86] $384^{2}$ 21.5 8.8 261 4.5 83.9 FastViT-MA36 [93] $256^{2}$ 42.7 7.9 314 4.7 83.9 FFNet-3 $256^{2}$ 48.3 10.1 490 4.5 83.9
Model Image Param FLOPs GPU Mobile Top-1 Size Throu. Latency Acc. (Px) (M) (G) (Img/Sec) (ms) (%) CoAtNet-2 [19] $224^{2}$ 75.0 15.7 - 24.4 84.1 HorNet-B7×7 [74] $224^{2}$ 87.0 15.6 268 8.5 84.2 EfficientNet-B7 [85] $600^{2}$ 66.3 37.0 182 45.0 84.3 ConvNeXt-L [60] $224^{2}$ 198.0 34.4 206 16.5 84.3 CMT-B [32] $256^{2}$ 45.7 9.3 253 13.2 84.5 EfficientViT-L1 [5] $224^{2}$ 52.7 5.3 354 82.8 84.5 Swin-B [59] $384^{2}$ 88.0 47.1 - 181.6 84.5 FFNet-3 $\mathbf{384^{2}}$ 48.3 22.8 360 9.1 84.5 MOAT-0 [107] $384^{2}$ 27.8 18.2 - 25.2 84.6 iFormer-S [81] $384^{2}$ 20.0 16.1 328 80.3 84.6 NFNet-F1 [4] $320^{2}$ 132.6 35.5 278 15.8 84.7 NFNet-F2 [4] $352^{2}$ 193.8 62.6 186 27.0 85.1 EfficientNetV2-M[86] $480^{2}$ 54.0 24.0 199 11.0 85.1 CoAtNet-1 [19] $384^{2}$ 42.2 27.4 - 41.6 85.1 ConvNeXt-B [60] $384^{2}$ 88.6 45.0 220 21.3 85.1 HorNet-B7×7 [74] $384^{2}$ 87.0 45.8 171 21.3 85.3 FFNet-4 $\mathbf{384^{2}}$ 79.2 43.1 264 15.2 85.3 ImageNet-1K distillation results SwiftFormer-L1 [78] $224^{2}$ 12.1 1.6 943 1.6 80.9 EfficientFormerV2-S2 [47] $224^{2}$ 12.6 1.3 671 2.0 81.6 FastViT-SA12 [93] $256^{2}$ 10.9 1.9 731 2.0 81.9 FFNet-1 $256^{2}$ 13.7 2.9 1113 1.8 82.1 SwiftFormer-L3 [78] $224^{2}$ 28.5 4.0 316 3.0 83.0 EfficientFormerV2-L [47] $224^{2}$ 26.1 2.6 542 2.8 83.3 FastViT-SA24 [93] $256^{2}$ 20.6 3.8 439 3.0 83.4 FFNet-2 $256^{2}$ 26.9 6.0 674 3.1 83.7 FFNet-3 $256^{2}$ 48.3 10.1 490 4.5 84.5
Architecture. Following convention, FFNet is structured into four stages, each consisting of multiple FFNet blocks for hierarchical feature extraction. In the token mixer, for efficiency, we adopt a kernel size of 3$\times$3 in the first two stages and switch to 7$\times$7 in later stages. In the channel mixer, we mainly use 7$\times$7 kernels, and the FFN’s expansion ratio is set to 3 by default ($d_{m}=3d$). The FFNet block is illustrated in Fig. 3 (c). For downsampling, we utilize a 7$\times$7 strided depthwise convolution followed by a 1$\times$1 convolution. We also use LayerScale [91] to make our deep models more stable during training. Model specifications are provided in Appendix G.1.
@@
Setup. For image classification, we evaluate FFNet on the ImageNet-1K [23] dataset. We follow the same training recipe in  [89, 110] for a fair comparison. Specifically, we train our models for 300 epochs using AdamW optimizer with an initial learning rate of 1×$\times 10^{-3}$ via cosine schedule and a weight decay of 0.05. We set the total batch size as 1024 and the input size as $256^{2}$. For $384^{2}$ resolution, we finetune the models for 30 epochs with weight decay 1×$\times 10^{-8}$, learning rate of 5×$\times 10^{-6}$, and batch size of 512. Following DeiT [89], we also report the results of hard distillation using RegNetY-16GF [73] (with top-1 accuracy of 82.9%) as the teacher model.
Table 2: COCO object detection and instance segmentation results using Cascade Mask R-CNN [6]. All backbone statistics are calculated with image size (1280, 800).
Backbone GPU Mobile $\text{AP}^{\text{box}}$ $\text{AP}^{\text{mask}}$ Throu. Latency Swin-T [59] 58 OOM 50.4 43.7 Swin-S [59] 34 OOM 51.9 45.0 ConvNeXt-T [60] 127 86.2 50.4 43.7 ConvNeXt-B [60] 54 173.1 52.7 45.6 HorNet-S7×7 [74] 64 124.9 52.7 45.6 HorNet-B7×7 [74] 44 182.2 53.3 46.1 FFNet-2 135 51.5 51.8 44.9 FFNet-3 90 80.5 52.8 45.6 FFNet-4 61 136.3 53.4 45.9
@@
Results. In Tab. 1, we compare our models against recent state-of-the-art models. The comparison results clearly show that our FFNet obtains the best accuracy-speed trade-off on two different testbeds (GPU and mobile device). Notably, our model demonstrates a significant efficiency margin compared to attention-based models [32, 59, 81, 5], particularly on the mobile device. Furthermore, FFNet outperforms CNN-based models such as ConvNeXt and HorNet, which use kernels of the same size, thereby affirming the superiority of the proposed MetaMixer framework. Our model also generalizes well to a larger image resolution and distillation-based training [89].
4.2 Object Detection and Instance Segmentation
@@
Setup. We evaluate the performance of FFNet as the backbone of Cascade Mask R-CNN [6] on the COCO object detection and instance segmentation tasks [52]. We adopt standard training recipe [59, 60] using 3$\times$ schedule with multi-scale training. All experiments are implemented on mmdetection [8] codebase.
@@
Results. The results presented in Tab. 2 indicate that FFNet models outperform baselines across all scales. Remarkably, our models demonstrate impressive efficiency improvements compared to the baselines with similar performance.
4.3 2D Semantic Segmentation
@@
Table 3: 2D semantic segmentation results. To facilitate a comprehensive evaluation of efficiency, we report backbone speed at the backbone-level and overall model speed at the system-level comparison. Backbone statistics are calculated with image size (2048, 512). The input size is (2048,1024) for Cityscapes. “SS”/“MS” denote single-scale and multi-scale testing, respectively. “*” indicates use of an RTX A6000 GPU.
Backbone Param FLOPs GPU Mobile mIoU Throu. Latency (M) (G) (Img/Sec) (ms) (SS/MS) Backbone-Level Comparison on ADE20K with UperNet 160K Swin-T [59] 60 945 60 OOM 44.5 / 45.8 Swin-B [59] 121 1188 28 OOM 48.1 / 49.7 ConvNeXt-T [60] 60 939 119 70 46.0 / 46.7 ConvNeXt-B [60] 122 1170 40 170 49.1 / 49.9 HorNet-S7×7 [74] 81 1030 57 112 49.2 / 49.8 HorNet-B7×7 [74] 121 1174 37 169 50.0 / 50.5 HorNet-BGF [74] 126 1171 - - 50.5 / 50.9 FFNet-2 58 942 135 45 47.1 / 47.8 FFNet-3 80 1010 90 69 49.6 / 50.2 FFNet-4 113 1158 60 124 50.7 / 51.7
Model Param FLOPs GPU* mIoU Latency (M) (G) (ms) (SS/MS) System-Level Comparison on ADE20K SETR-MLA [120] 310 368 82 47.5 / 49.4 Segformer-B3 [106] 47 79 256 49.4 / 50.0 MaskFormer [12] 63 79 - 49.8 / 51.0 SegNeXt-B [33] 28 35 127 48.5 / 49.9 EfficientViT-L1 [5] 40 36 80 49.2 / - $\text{FFNet}_{\textit{seg}}$ 68 74 65 50.1 / 51.2 System-Level Comparison on Cityscapes Segformer-B3 [106] 47 963 252 81.7 / 83.3 SegNeXt-L [33] 49 578 307 83.2 / 83.9 $\text{FFNet}_{\textit{seg}}$ 68 577 224 83.2 / 84.1
@@
Architecture. For comparison at the backbone level, we utilize ImageNet-pretrained models as the backbone of UperNet [105]. Furthermore, to enable a system-level comparison, we propose $\text{FFNet}_{\textit{seg}}$, a model similar to the original FFNet but incorporating task-specific modifications. $\text{FFNet}_{\textit{seg}}$ adopts the standard backbone-head (encoder-decoder) architecture. Similar to the backbone, the head also employs FFNet blocks, thereby avoiding complex operations such as global self-attention [12] and matrix decomposition [33]. Following  [33], we input the features from the last three stages into the head, where they are concatenated and processed through an FFN. Detailed configurations of the head for each datset will be provided in our official GitHub repository.
Recent works [5, 33, 118] demonstrate that long-range interactions and multi-scale features are crucial for competitive performance in segmentation models. Accordingly, to enlarge the receptive field, we employ 9$\times$9 kernels in the final two stages of the backbone and the head. Additionally, to effectively process features of objects with varying sizes, we combine small-kernel and strip convolutions [38, 71] (9$\times$1, 1$\times$9) with large-kernel convolution, enhancing flexible feature extraction. Further details and a visual schematic can be found in Appendix F and Tab. 10.
Setup. We evaluate the effectiveness of our FFNet on two representative benchmark datasets: ADE20K [121] and Cityscapes [16]. For backbone-level comparison, our models are trained for 160K-iterations with a batch size of 16, following standard recipe [59, 60]. For system-level comparisons, we adopt the same training settings in SegNeXt [33] and use ImageNet-1K pretrained models for initialization. We also use MMSegmentation toolbox [14] for all semantic segmentation experiments.
Results. Tab. 3 shows that our methods excel not only as backbones but also as standalone segmentors, highlighting our model’s superior performance-speed trade-off. For example, FFNet-4 achieves a $+1.8$ mIoU gain over ConvNeXt-B, while providing 50% improvement in GPU speed and 37% increase in mobile speed. Also, $\text{FFNet}_{\textit{seg}}$ consistently outperforms task-specific prior arts.
4.4 Super-Resolution
Architecture. $\text{FFNet}_{\textit{sr}}$ consists of three parts: shallow feature extractor (HSF), deep feature extractor (HDF), and upscale module (HRec) following residual-in-residual architecture [115]. HSF extracts shallow features from a Low-Resolution (LR) image using a 3$\times$3 convolution. The deep feature extractor HDF, comprising a stack of FFNet blocks, refines shallow features iteratively. Specifically, $\text{FFNet}_{\textit{sr}}$-light and $\text{FFNet}_{\textit{sr}}$ employ 36 and 48 blocks, respectively. We use 3$\times$3 depthwise convolutions across all variants, with 96 base channels and an FFN expansion ratio of 2. In line with findings from EDSR [51], we also observe similar performance degradation when using Batch Normalization (BN) and consequently exclude BN from the FFNet Block. We use anchor-based long residual connection [28, 125] where the LR image is repeated RGB-wise and added at the end of the HDF. Finally, HRec reconstructs the High-Resolution image by up-sampling refined features (see Appendix G.2 for more details).
Setup. $\text{FFNet}_{\textit{sr}}$-light and $\text{FFNet}_{\textit{sr}}$ is evaluated on the DIV2K [51] and DF2K (DIV2K + Flicker2K [87]) dataset, respectively. Five standard datasets are adopted for testing: Set5 [3], Set14 [113], BSD100 [63], Urban100 [39], and Manga109 [64]. We use PSNR and SSIM to evaluate the SR performance on the Y channel from the YCbCr color space after cropping 2 pixels at boundary [41]. In each training batch, we randomly crop LR patches of size 96$\times$96 as input and augment the patches with random horizontal flip and rotations. Our models are trained by minimizing L1 loss and frequency reconstruction loss [92, 84] with a weight factor 0.02. We use Adam optimizer for training with a batch size 64 for 500K iterations. The initial learning rate is set to 0.001 and updated by the cosine annealing scheme. All latency and memory usage are measured corresponding to an HR image of size 1280$\times$720.
Results. In Tab. 4, we compare our models against recent state-of-the-art SR models. Notably, $\text{FFNet}_{\textit{sr}}$ significantly outperforms SR specialists in efficiency, operating seamlessly even on resource-constrained devices, demonstrating their feasibility for real-world applications. $\text{FFNet}_{\textit{sr}}$, employing 3$\times$3 convolutions within the MetaMixer framework, not only captures high-frequency details effectively but also allows for the stacking of more blocks within the same computational budget compared to the baselines [84, 11, 95] that use complex mixers, thereby enhancing its ability to capture long-range contexts. The impressive results suggest that our mixer design is not only effective for high-level tasks such as recognition but also well-suited for solving low-level problems. Visual results of the $\text{FFNet}_{\textit{sr}}$ are provided in Appendix I.
Table 4: Comparisons on representative benchmark datasets for Super-Resolution networks at $\times$2 scale. All results are calculated on the Y-channel. All statistics are calculated corresponding to an HR image of size 1280$\times$720. Memory denotes maximum GPU memory consumption during the inference phase. Best and second-best results are bolded and underlined, respectively. $\dagger$ denotes we re-implement model since their codes are unavailable.

### v2

@@ 3.3 FFNet as a General Backbone
Given our proposed FFNified attention and off-the-shelf ConvNeXt block as primary building mixers, we introduce Fast-Forward Network (FFNet), a new network family that achieves competitive results compared to sophisticated methods on diverse benchmarks, utilizing extremely simple operators based on MetaMixer. The basic block of our FFNet is illustrated in Fig. 3 (c). Moreover, our model consists solely of hardware-friendly operations, ensuring high efficiency across various devices (see Tables 1 and S.1).
@@
4 Experiments
To validate the inherent effectiveness of the MetaMixer framework and its resulting convolutional mixer design, we conduct a comprehensive comparison of our Fast-Forward Network (FFNet) with state-of-the-art baselines across a range of tasks, such as image classification, object detection, 2D instance/semantic segmentation, super-resolution, 3D semantic segmentation, and long-term time series forecasting. Lastly, building upon MetaMixer’s holistic view of representation learning, we harmoniously integrate attention and FFNified attention to improve the performance of ImageNet generation models.
@@ 4 Experiments
Speed Measurement. Following the same protocol as [94, 49], mobile latency is measured on iPhone 12 (iOS 16.5) using models exported by CoreML tools, while GPU throughput is measured with an NVIDIA RTX 4090 GPU. For GPU throughput, we compile the traced model with TensorRT (v8.6.1.post1). Unless otherwise specified, all methods are evaluated using their respective input sizes and a batch size of 1, following the aforementioned procedure.
Model Image Param FLOPs GPU Mobile Top-1
Size Throu. Latency Acc.
(Px) (M) (G)
(imgs/s)
(ms) (%)
EfficientNet-B1 [85] $240^{2}$ 7.8 0.7 1115 1.8 79.1
FastViT-SA12 [94] $256^{2}$ 10.9 1.9 1274 2.0 80.6
MPViT-XS [47] $224^{2}$ 11.0 2.9 504 45.2 80.9
Swin-T [58] $224^{2}$ 29.0 4.5 - 24.8 81.3
FasterNet-S [8] $224^{2}$ 31.1 4.6 1040 2.5 81.3
FFNet-1 $256^{2}$ 13.7 2.9 1430 1.8 81.3
ConvNeXt-T [59] $224^{2}$ 29.0 4.5 952 3.7 82.1
SMT-T [54] $224^{2}$ 11.5 2.4 802 23.2 82.2
RMT-T [29] $224^{2}$ 14.0 2.5 399 - 82.4
FastViT-SA24 [94] $256^{2}$ 20.6 3.8 596 3.0 82.6
HorNet-T7×7 [74] $224^{2}$ 22.0 4.0 445 3.4 82.8
EfficientNet-B4 [85] $380^{2}$ 19.3 4.2 652 5.2 82.9
Swin-S [58] $224^{2}$ 50.0 8.7 - 29.3 83.0
FFNet-2 $256^{2}$ 26.9 6.0 776 3.1 82.9
MOAT-0 [106] $224^{2}$ 27.8 5.7 - 5.1 83.3
iFormer-S [81] $224^{2}$ 20.0 4.8 500 27.3 83.4
MogaNet-S [48] $224^{2}$ 25.0 5.0 417 4.3 83.4
Swin-B [58] $224^{2}$ 87.8 15.4 - 33.5 83.5
CMT-S [33] $224^{2}$ 25.1 4.0 407 11.3 83.5
FasterNet-L [8] $224^{2}$ 93.5 15.5 498 6.8 83.5
EfficientNet-B5 [85] $456^{2}$ 30.0 9.9 439 11.0 83.6
SMT-S [54] $224^{2}$ 20.5 4.7 421 41.5 83.7
HorNet-S7×7 [74] $224^{2}$ 49.5 8.8 351 5.4 83.8
ConvNeXt-B [59] $224^{2}$ 88.6 15.4 399 8.1 83.8
BiFormer-S [125] $224^{2}$ 26.0 4.5 162 - 83.8
EfficientNetV2-S[86] $384^{2}$ 21.5 8.8 298 4.5 83.9
FocalNet-B (LRF) [107] $224^{2}$ 88.7 15.4 287 10.0 83.9
FastViT-MA36 [94] $256^{2}$ 42.7 7.9 422 4.7 83.9
FFNet-3 $256^{2}$ 48.3 10.1 540 4.5 83.9
RMT-S [29] $224^{2}$ 27.0 4.5 239 - 84.1
DaViT-S [23] $224^{2}$ 49.7 8.8 177 16.4 84.2
HorNet-B7×7 [74] $224^{2}$ 87.0 15.6 280 8.5 84.2
ConvNeXt-L [59] $224^{2}$ 198.0 34.4 218 16.5 84.3
MogaNet-B [48] $224^{2}$ 44.0 9.9 222 8.2 84.3
CMT-B [33] $256^{2}$ 45.7 9.3 294 13.2 84.5
Swin-B [58] $384^{2}$ 88.0 47.1 - 181.6 84.5
MaxViT-S [93] $224^{2}$ 69.0 11.7 141 188.1 84.5
FFNet-3 $384^{2}$ 48.3 22.8 387 9.1 84.5
MOAT-0 [106] $384^{2}$ 27.8 18.2 - 25.2 84.6
iFormer-S [81] $384^{2}$ 20.0 16.1 334 80.3 84.6
FasterViT-3 [36] $224^{2}$ 159.3 18.2 - 60.8 84.9
RMT-B [29] $224^{2}$ 54.0 9.7 162 - 85.0
NFNet-F2 [4] $352^{2}$ 193.8 62.6 194 27.0 85.1
MogaNet-XL [48] $224^{2}$ 181.0 34.5 118 22.3 85.1
ConvNeXt-B [59] $384^{2}$ 88.6 45.0 229 21.3 85.1
MaxViT-T [93] $384^{2}$ 31.0 17.7 149 247.6 85.2
HorNet-B7×7 [74] $384^{2}$ 87.0 45.8 175 21.3 85.3
FFNet-4 $384^{2}$ 79.2 43.1 282 15.2 85.3
ImageNet-1k distillation results
SwiftFormer-L1 [78] $224^{2}$ 12.1 1.6 943 1.6 80.9
EfficientFormerV2-S2
 [50]
$224^{2}$ 12.6 1.3 761 2.0 81.6
FastViT-SA12 [94] $256^{2}$ 10.9 1.9 1274 2.0 81.9
FFNet-1 $256^{2}$ 13.7 2.9 1430 1.8 82.1
SwiftFormer-L3 [78] $224^{2}$ 28.5 4.0 609 3.0 83.0
EfficientFormerV2-L [50] $224^{2}$ 26.1 2.6 598 2.8 83.3
FastViT-SA24 [94] $256^{2}$ 20.6 3.8 596 3.0 83.4
FFNet-2 $256^{2}$ 26.9 6.0 776 3.1 83.7
FFNet-3 $256^{2}$ 48.3 10.1 540 4.5 84.5
@@
Table 1: ImageNet-1k classification. “-” denotes that model could not be reliably exported by either TensorRT or Core ML Tools.
4.1 FFNet for Efficient Image Recognition
@@
ImageNet Classification. Following convention, FFNet is structured into four stages, each consisting of multiple FFNet blocks for hierarchical feature extraction. The FFNet block is illustrated in Fig. 3 (c). In the token mixer, for efficiency, we adopt a kernel size of 3$\times$3 in the first two stages and switch to 7$\times$7 in later stages. In the channel mixer, we mainly use 7$\times$7 kernels, and the expansion ratio is set to 3 by default. We evaluate FFNet on the ImageNet-1k [22] dataset. We follow the same training recipe in  [90, 109] for a fair comparison. Specifically, we train our models for 300 epochs using AdamW optimizer with an initial learning rate of 1×$\times 10^{-3}$ via cosine schedule and a weight decay of 0.05. We set the total batch size as 1024 and the input size as $256^{2}$. For $384^{2}$ resolution, we finetune the models for 30 epochs with weight decay 1×$\times 10^{-8}$, learning rate of 5×$\times 10^{-6}$, and batch size of 512. Following DeiT [90], we also report the results of hard distillation using RegNetY-16GF [73] as the teacher model. Further details on the model specifications are provided in Appendix F.1.
@@
Results. In Tab. 1, we compare our models against recent state-of-the-art methods. Notably, the comparative results clearly show that our FFNet achieves the best speed-accuracy trade-off on two testbeds (GPU and mobile device). In particular, our model demonstrates a substantial efficiency improvement over attention-based models [33, 58, 81, 93, 125, 23, 29, 54], especially on the mobile device. Furthermore, FFNet outperforms CNN-based models such as ConvNeXt, HorNet, and MogaNet, which use kernels of the same size, thereby affirming the superiority of the proposed MetaMixer framework. FFNet also generalizes well to a larger image resolution and distillation-based training [90]. A thorough comparison with relevant methods is provided in Appendix C and Tab. 8.
@@
Object Detection and Instance Segmentation. We evaluate the performance of FFNet as the backbone of Cascade Mask R-CNN [6] on the COCO instance segmentation task [53]. We adopt standard training recipe [58, 59] using 3$\times$ schedule with multi-scale training. All experiments are implemented on MMDetection [9] codebase.
@@
Results. The results presented in Tab. 3 indicate that FFNet models outperform baselines across all scales. Remarkably, our models demonstrate impressive efficiency gains compared to the SOTA baselines. For example, at comparable performance levels, FFNets run nearly twice as fast as RMT-S on the GPU and ConvNeXt-B on the iPhone 12.
Backbone GPU Mobile $\text{AP}^{\text{box}}$ $\text{AP}^{\text{mask}}$
[…]
SMT-S [54] 29 OOM 51.9 44.7
Swin-S [58] 34 OOM 51.9 45.0
ConvNeXt-T [59] 139 86.2 50.4 43.7
ConvNeXt-B [59] 54 173.1 52.7 45.6
HorNet-S7×7 [74] 66 124.9 52.7 45.6
HorNet-B7×7 [74] 44 182.2 53.3 46.1
RMT-S [29] 32 - 53.2 46.1
FFNet-2 141 51.5 51.8 44.9
FFNet-3 92 80.5 52.8 45.6
FFNet-4 62 136.3 53.4 45.9
Table 2: COCO object detection and instance segmentation. All backbone statistics are calculated with image size (1280, 800).
Method Param FLOPs GPU Mobile mIoU
Latency Latency SS / MS
(M) (G) (ms) $\downarrow$ (ms) $\downarrow$ (%)
Backbone-level Comparison on ADE20K with UperNet 160K
Swin-B [58] 121 1188 36 OOM 48.1 / 49.7
ConvNeXt-T [59] 60 939 8 70 46.0 / 46.7
ConvNeXt-B [59] 122 1170 23 170 49.1 / 49.9
SMT-S [54] 50 935 35 1198 49.2 / 50.2
DaViT-B [23] 121 1175 24 OOM 49.4 /   -
RMT-S [29] 56 937 34 - 49.8 /   -
HorNet-S7×7 [74] 81 1030 17 112 49.2 / 49.8
HorNet-B7×7 [74] 121 1174 26 169 50.0 / 50.5
HorNet-BGF [74] 126 1171 - - 50.5 / 50.9
MogaNet-B [48] 74 1050 30 310 50.1 /   -
FocalNet-B (LRF) [107] 126 1192 23 - 50.5 / 51.4
FFNet-2 58 942 7 45 47.1 / 47.8
FFNet-3 80 1010 11 69 49.6 / 50.2
FFNet-4 113 1158 16 124 50.7 / 51.7
System-level Comparison on ADE20K
SETR-MLA [119] 310 368 82 - 47.5 / 49.4
Segformer-B3 [105] 47 79 256 - 49.4 / 50.0
MaskFormer [13] 63 79 - - 49.8 / 51.0
SegNeXt-B [34] 28 35 127 - 48.5 / 49.9
EfficientViT-L1 [5] 40 36 80 - 49.2 /   -
$\text{FFNet}_{\textit{seg}}$ 68 74 65 - 50.1 / 51.2
System-level Comparison on Cityscapes
Segformer-B3 [105] 47 963 252 - 81.7 / 83.3
SegNeXt-L [34] 49 578 307 - 83.2 / 83.9
$\text{FFNet}_{\textit{seg}}$ 68 577 224 - 83.2 / 84.1
@@
Table 3: 2D semantic segmentation. For a thorough assessment of efficiency, we report backbone latency at the backbone level and total forward pipeline latency at the system level comparison. System-level latency is exceptionally measured on an RTX A6000 with PyTorch. Backbone statistics are calculated with image size (2048, 512). For Cityscapes, the input size is (2048, 1024). “SS”/“MS” denote single- and multi-scale testing, respectively.
@@
2D Semantic Segmentation. For comparison at the backbone level, we utilize ImageNet-pretrained models as the backbone of UperNet [104]. Furthermore, to enable a system-level comparison, we propose $\text{FFNet}_{\textit{seg}}$, a model similar to the original FFNet but incorporating task-specific modifications. $\text{FFNet}_{\textit{seg}}$ adopts the standard backbone-head (encoder-decoder) architecture. Similar to the backbone, the head also employs FFNet blocks, thereby avoiding complex operations such as global self-attention [13] and matrix decomposition [34]. Following  [34], we input the features from the last three stages into the head, where they are concatenated and processed through an FFN. More details are presented in Appendix F.1.